If the temperature is 104 oF and the relative humidity is 82.3 %, what is the humidity? Record your answer to three decimal places
1
Expert's answer
2012-09-27T10:55:37-0400
Tc = 5.0 / 9.0 * (Tf - 32.0) = 5.0 / 9.0 * (104.0 - 32.0) = 40 *C
According
to the table, humidity is 42.070159722 = 42.070